OpenStudy (anonymous):

A recipe calls for oz of water 1/2 c. Of sugar 1 oz. Of lemon juice and 4 mint leaves how much liquid does a triple recipe call for

OpenStudy (chriss):

How many ounces of water?

OpenStudy (anonymous):

10 sorry

OpenStudy (chriss):

Ok, so to triple it, you multiply 3 times the amount of the liquid ingredients. The liquid ingredients are water and lemon juice of which you have 10 oz and 1 oz respectively. So 3*10 = 30 and 3*1 = 3 and 30+3=33 So you have 33 oz of liquid in a triple recipe.